# Maximal Extractable Value in Decentralized Finance: Taxonomy, Detection, and Mitigation

### v1

@@ 4.3 Article Selection
This step involves selecting articles for inclusion in the survey. Figure 1 illustrates the flowchart of the article selection process, following the Preferred Reporting Items for Systematic Reviews and Meta-Analyses (PRISMA) 2020 guidelines [17]. As depicted, a total of 215 records were retrieved through the database searches, along with 7 records from the manual search. Initially, 77 duplicate records were removed, leaving 145 records for the screening stage. These records were screened based on their titles and abstracts, resulting in the exclusion of 91 records, which reduced the total number of articles to 54. A full-text analysis of these 54 articles was then conducted to identify the most relevant ones for inclusion in the survey. Applying the exclusion criteria, 20 articles were excluded, leading to the inclusion of 34 articles in this survey.
@@ 5 Taxonomy of MEV Transactions
MEV transactions can be classified into several types based on the transaction details (such as token transfer and add/remove liquidity) and their order within a block relative to a victim transaction888A transaction in the public mempool that presents an opportunity for a searcher to generate profits.. Figure 2 presents our taxonomy of different MEV transaction types. These transactions can either be value-diverting (front-running, back-running, sandwich, and time-bandit) or both value-creating and value-diverting (arbitrage and liquidation). This section provides a detailed explanation of each transaction type, including real-world examples where applicable.
@@ 5.2 Back-running
In back-running, a searcher monitors the mempool for a victim transaction, causing price slippage. The searcher then submits a back-running transaction after the victim transaction at a lower gas price. Back-running transactions could lead to market inefficiency by further increasing price slippage. Furthermore, it reduces profitable opportunities for other users due to liquidity drain. Figure 5 presents an example of a back-running transaction.
@@ 5.5 Liquidation
One of the most prominent applications of DeFi is lending and borrowing assets. Borrowers can obtain loans from lending pools, such as Aave999https://aave.com/ and Compound101010https://compound.finance/, by providing collateral as security. Lenders earn profits through the interest paid by borrowers. To secure lenders, the value of the collateral must be more than a certain threshold (for instance, 150% of the loan). However, due to market volatility, the collateral value may fall below this threshold. Lending protocols allow any user to liquidate loans when the collateral value approaches the threshold. The borrower may prevent liquidation by performing a trade that increases the collateral’s value. Alternatively, a non-borrower user can initiate a liquidation transaction to repay the debt and acquire the collateral at a lower/discounted price, thereby protecting the lender from potential losses. This type of liquidation transaction is a value-creating MEV transaction. On the other hand, searchers may exploit these liquidation opportunities to extract MEV by performing value-diverting transactions either in one of the following ways: 1) a searcher monitors the mempool and front-runs the transaction of a competing liquidator to capture their profit, 2) a searcher monitors transactions that would create the liquidation opportunity (referred to as forced liquidation) and then back-runs the transaction to acquire the collateral [28], 3) a searcher performs a forced liquidation, by front-running a transaction to create a liquidation opportunity and then back-running another transaction to claim the collateral. Furthermore, a collusion of searchers could manipulate market prices to extract MEV from multiple blocks by performing liquidation transactions [29]. These types of liquidation transactions have a detrimental impact on the borrowers and the market as it increases collateral requirements. Figure 18 illustrates an example of a liquidation transaction. Liquidations are of two types: fixed spread-based and auction-based [28]
@@ 5.6 Time-bandit
The previously described MEV transactions profit from re-arranging, inserting, or removing transactions in a block that is to be mined. On the contrary, time-bandit MEV transactions aim to extract MEV from already mined block [11]. The searcher, on spotting an MEV opportunity in a finalized block, would fork the blockchain by re-mining a block with MEV-acquiring transactions. Time-bandit transactions are not limited to MEV from DEXes [11]. Time-bandit MEV leads to blockchain instability due to re-mining of blocks.
@@ 6.1 Front-running Detection
Torres et al. [3] focused on the detection of displacement and suppression front-running transactions. The displacement detection approach begins by identifying the front-running () and victim () transaction pairs using a sliding window that iterates over each transaction within blocks included in the window. Each transaction input bytes is divided into 4-byte n-grams. If 95%111111100% matching criteria is not considered as might be encapsulated within along with some metadata. of n-grams in a transaction corresponds to a previous transaction, a match is reported. For a matched transaction pair, the proposed heuristic then determines if is a displacement transaction if and have different senders as well as receivers (different receivers ensure that bot contracts perform displacement transactions), the gas price of the is larger than that of , and ratio between byte sequences of and is at least 25%. The proposed approach is only limited to displacement transactions performed by bot contracts. Furthermore, it may fail to detect displacement if and do not occur within the same sliding window, leading to false negatives.
@@ 6.2 Sandwich Detection
Torres et al. [3] proposed a heuristic to detect single DEX and cross-DEX sandwich transactions. The heuristic begins by identifying all possible transfer events121212A transfer event (containing sender’s address, receiver’s address, and traded amount) is emitted whenever a token is traded. within a block. The heuristic flags the presence of sandwich transactions if three events (i.e., front-run (), victim (), and back-run ()) satisfy the following conditions: Sender of and is the receiver of , and receiver of is the sender of , difference between token value of and of is less than or equal to 1%, token contract addresses for three events are same, transaction hashes for the events are different, transaction index of is smaller than that of which in turn is smaller than that of , and gas price of is greater than that of which in turn is greater than or equal to that of . The same heuristic is employed by Weintraub et al. [28] to detect sandwich transactions executed via Flashbots. However, this heuristic fails to identify sandwich transactions where the difference between and exceeds 1%131313Front-run transaction index:7, victim transaction index:9, back-run transaction index:10 (Block 20633424).
To detect similar types of sandwich transactions, Qin et al [15] proposed a heuristic based on the following conditions: front-run (), victim (), and back-run () transactions are in the same block, and this order, should map to one and only one , swap directions of and are the same, and that of is the opposite, either and have the same sender or same receiver smart contract, and difference between token value swapped in and swapped back in is within 10% bound. However, the heuristic cannot detect sandwich attacks where the difference between and exceeds 10%141414Front-run transaction index:21, victim transaction index:22, back-run transaction index:23 (Block 20619305).
The heuristics employed by [3, 28, 15] fail to identify sandwich transactions where the indexes of , , are not consecutive151515Front-run transaction index:0, victim transaction index:2, back-run transaction index:3 (Block 20619280) or multiple sandwich transactions occur within a single block. Furthermore, they are not capable of detecting burger, dagwood, and liquidation sandwich transactions. These limitations result in a high rate of false negatives.
@@
The previously discussed heuristics for detecting sandwich transactions [3, 28, 15, 23, 25] rely on decoding smart contracts to identify function calls and transfer events. As a result, these methods are limited to smart contracts with accessible binary files for decoding. To address this gap, Park et al. [20] proposed a heuristic for sandwich detection that focuses on token transfer data instead of contract data. The proposed heuristic first identifies the pairs of transactions ( and ) sharing the same recipient address within a single block. Then, the transaction pairs are filtered out if one of the following conditions is met: only one token is transferred in either of the transactions, if profits for all token transfers in are greater than zero, if have all trades with zero profits, if set of token contract address in and are not same, or total profit from and is less than zero. Compared to Flashhbots with a 95.15% F1-score, the proposed approach achieved an F1-score of 98.83%. However, since this approach relies on token transfer data, identifying specific function calls becomes difficult, resulting in false positives. Furthermore, the approach is unable to detect burger, dagwood, and liquidation sandwich transactions.
@@ 6.3 Arbitrage Detection
Daian et al. [11] conducted a preliminary quantification of arbitrage in Ethereum based on a manual analysis of atomic arbitrage transactions. Later, Qin et al. [15] proposed a heuristic to detect arbitrages. According to this heuristic, a transaction is considered arbitrage if the following conditions are met: the transaction contains more than one swap event, and all events are confined within a single transaction, all swap events must form a loop, with the input of one swap being the output of the previous swap, and the input of the first swap being the output of the last swap, and the input value of each swap event should be less than or equal to the output value of the previous event. This heuristic was also employed by Weintraub et al. [28] to detect arbitrage transactions through Flashbots and Flash Loans. However, the heuristic has several limitations. It fails to identify a cyclic arbitrage transaction if the value of the input token for a swap exceeds the value of the output token from the previous swap 161616Transaction index:2 (Block 14628933). Furthermore, it cannot detect reverse-ordered cyclic, unordered cyclic, cyclic NFT, or cyclic multi-address arbitrage transactions. Additionally, because it only considers transactions that form a cycle of swap events, it is unable to identify burn and mint or set token arbitrage transactions.
@@
Later, Park et al. [20] introduced a heuristic for detecting cyclic NFT arbitrage transactions involving swaps between ERC-20 tokens and ERC-721 NFTs. They also proposed a graph neural network model to identify ordered cyclic, cyclic multi-address, burn and mint, and set token arbitrage transactions. The model uses features related to acquired profit, the number of tokens sent and received, involved addresses, and the number of transfers sent and received to construct a graph. The authors implemented Graph Convolutional Networks (GCN), GraphSAGE, and Graph Attention Networks (GAT). While their approach achieved high F1-scores of 99.54% and 97.59% for training and test data, respectively, compared to Flashbots’ 68.65% and 62.85%, it may suffer from a high false positive rate due to the dynamic nature of the DEX market. Furthermore, a decrease in the performance of test data compared to train data could be due to overfitting and requires further analysis. The approach fails to detect reverse-ordered and unordered cyclic arbitrage transactions.
@@ 6.4 Liquidation Detection
Weintraub et al. [28] developed a method to identify liquidation transactions via Flashbots and Flash loans by searching for LiquidationCall and LiquidateBorrow events on the Aave and Compound lending platforms, respectively. Furthermore, Qin et al. [15] extended the scope by detecting liquidation events across Aave, Compound, and dYdX platforms. However, these approaches are limited to specific platforms.
@@ 6.5 Multi-type MEV Detection
Varun et al. [30] proposed a neural network-based approach to detect front-running and sandwich transactions, specifically targeting displacement and suppression front-running and single and cross-DEX sandwich transactions. Their approach involves extracting features for a transaction based on prior transactions and inputting them into a pre-trained Multi-Layer Perceptron (MLP) model to identify MEV transactions. The features extracted include the transaction’s gas price, the mean and standard deviation of transactions’ gas prices over the last 10 blocks, the mean and standard deviation of the same account’s historical gas prices, gas token usage, and a predicted gas price. The predicted gas price is obtained using a Long Short-Term Memory (LSTM) model based on the gas prices of the previous 12 non-MEV transactions. Using the MEV transactions data identified by Torres et al. [3] for training and validation, the approach achieved accuracies of 85.53% for suppression front-running transactions, 88.58% for displacement front-running transactions, and 89.08% for sandwich transactions. However, the proposed learning-based model may be prone to overfitting as it depends on a tailored dataset and might fail to generalize to a wider range of front-running and sandwich transactions.
@@
Li et al. [21] introduced a learning-based approach to detect sandwich, arbitrage, and liquidation MEV transactions. The authors proposed ACTLIFTER, which identifies MEV transactions within Flashbot bundles by mapping them to asset transfer patterns. Additionally, they developed ACTCLUSTER to discover new categories of MEV transactions within the sandwich, arbitrage, and liquidation types. ACTCLUSTER clusters MEV actions identified by ACTLIFTER by converting the bundle into a matrix where each DeFi action is represented as an action block. This matrix is processed using a Convolutional Neural Network (CNN) to extract features, which are then flattened and fed into three fully connected neural network layers. MEV actions are initially labeled as a sandwich, cyclic arbitrage, or liquidation transaction, with these labels being extended after each clustering round using DBSCAN. ACTLIFTER achieved better detection results compared to Etherscan171717https://etherscan.io/, accessed on 05 September 2024 and DeFiRanger [31]. However, ACTLIFTER is limited to identifying swap activities involving a token pair and fails to detect token chains.
@@
Table III presents the capabilities of the proposed detection approaches in detecting different types of MEV transactions. In the context of front-running MEV transactions, no approach focuses on replacement transactions. Regarding sandwich MEV transactions, most works focus on detecting single and cross-DEX sandwich transactions [3, 28, 15, 23, 20, 30, 21]. Only a few focus on detection of burger [23, 21], dagwood [23], and liquidation sandwich [25, 21] transactions. No single approach is capable of detecting all types of sandwich transactions. Furthermore, all these approaches are limited to single-block sandwich transactions. Consequently, the detected sandwich transactions represent only lower bounds and could result in a high number of false negatives [3, 15]. For arbitrage detection, most works focus on ordered cyclic transactions [28, 15, 23, 20, 27, 21]. Additionally, most of the works [28, 15, 23, 20, 21] are limited to atomic arbitrage transactions, where multiple swap events occur within a single transaction. Only [27] addresses arbitrages performed through both atomic and non-atomic transactions. Lastly, only [21] focuses on the detection of liquidation MEV transactions.
@@
To summarize, there is no work that is capable of detecting all types of MEV transactions. Furthermore, the detection results presented by the studies discussed above [3, 28, 15, 23, 25, 20, 11, 27, 30, 21] are only estimates of MEV, as there is no definitive ground truth for MEV [15]. In addition, these works represent reactive MEV detection approaches, focusing on quantifying MEV acquired in the past. Estimating potential or future MEV via a proactive approach is challenging because the MEV obtained by each searcher varies based on the optimization techniques used to maximize MEV and the targeted victim transactions [16].
| MEV transactions | ||||||||||||||||||
| Front-running | Sandwich | Arbitrage | ||||||||||||||||
@@
| Detection approach | Displa- cement | Repla- cement | Suppre- ssion | Back- running | Single DEX | Cross- DEX | Burger | Dagwood | Liquid- ation | Ordered cyclic | Reverse- ordered cyclic | Unordered cyclic | Cyclic NFT | Cyclic multi- address | Burn and mint | Set token | Liquid- ation | Time- bandit |
| [3] | ||||||||||||||||||
| [28] | ||||||||||||||||||
| [15] | ||||||||||||||||||
| [23] | ||||||||||||||||||
| [25] | ||||||||||||||||||
| [20] | ||||||||||||||||||
| [27] | ||||||||||||||||||
| [30] | ||||||||||||||||||
| [21] | ||||||||||||||||||
@@
| : detected; : not detected; : partially detected; 1if and occur within the same block range window; 2if multiple suppression transactions are sent to the same bot contract; 3if difference between and is 1%; 4if the input value of a swap event the output value of previous event; 5if difference between and is 10%; 6if and has the same swap directions | ||||||||||||||||||

### v2

@@ 4.3 Article Selection
This step involves selecting articles for inclusion in the survey. We followed the Preferred Reporting Items for Systematic Reviews and Meta-Analyses (PRISMA) 2020 guidelines [20] for article selection. A total of 215 records were retrieved through the database searches, along with 21 records from the manual search. Initially, 77 duplicate records were removed, leaving 159 records for the screening stage. These records were screened based on their titles and abstracts, resulting in the exclusion of 91 records, which reduced the total number of articles to 68. A full-text analysis of these 68 articles was then conducted to identify the most relevant ones for inclusion in the survey. Applying the exclusion criteria, 21 articles were excluded, leading to the inclusion of 47 articles in this survey
@@ 5 Taxonomy of MEV Transactions
MEV transactions can be classified into several types based on the transaction details (such as token transfer and add/remove liquidity) and/or their order within a block relative to a victim transaction. Figure 1 presents our taxonomy of MEV transaction types and compares them across various metrics. In addition, the figure depicts which DeFi applications are affected by each MEV transaction type. The analyses for gas price sensitivity and network congestion assume MEV transactions are executed by a single searcher without PGAs. The MEV volume statistics are obtained from EigenPhi999https://eigenphi.io/, accessed on 10 February 2025. This section provides a detailed explanation of each transaction type, including real-world examples where applicable.
@@ 5.2 Back-running
In back-running, a searcher monitors the mempool for a victim transaction that causes a price discrepancy, which can be arbitraged. The searcher then submits a back-running transaction after the victim transaction, at a lower gas price, to capture the price difference. Back-running can be value-creating by reducing price discrepancies and contributing to market stability. However, it reduces profitable opportunities (value-diverting) for subsequent users due to liquidity drain, particularly for those who execute opposite-direction trades. Figure 4 presents an example of a back-running transaction.
@@ 5.4.7 Set token
5.4.8 Cross-domain
A cross-domain arbitrage transaction exploits price discrepancies between asset pairs across two or more domains, such as Layer 1, Layer 2, side-chains, and CEXes [31]. When this arbitrage occurs between an on-chain DEX and an off-chain CEX, it is known as DEX-CEX arbitrage. DEX-CEX arbitrage opportunities arise due to price volatility on CEXes, which can create temporary deviations from prices on on-chain DEXes [32]. If cross-domain arbitrage occurs between DEXes across different blockchains (e.g., Ethereum and BNB chain), it is referred to as cross-chain arbitrage [33]. When this arbitrage occurs between different execution layers of the same blockchain (e.g., Layer 1 and Layer 2) or between Layer 2 solutions or side-chains, it is referred to as cross-layer arbitrage. When cross-domain arbitrage occurs between different rollups on the same base-layer blockchain (e.g., Arbitrum and Optimism), it is referred to as cross-rollup arbitrage [34]. Cross-domain arbitrage transactions are non-atomic since different domains do not share a unified execution environment. This increases the risk for searchers, as not all transactions may execute successfully, potentially affecting profitability.
@@ 5.5 Liquidation
Lending protocols allow any user to liquidate loans when the collateral value approaches the threshold. The borrower may prevent liquidation by performing a trade that increases the collateral’s value. Alternatively, a non-borrower user can initiate a liquidation transaction to repay the debt and acquire the collateral at a lower/discounted price, thereby protecting the lender from potential losses. This type of liquidation transaction is a value-creating MEV transaction. On the other hand, searchers may exploit these liquidation opportunities to extract MEV by performing value-diverting transactions either in one of the following ways: 1) a searcher monitors the mempool and front-runs the transaction of a competing liquidator to capture their profit, 2) a searcher monitors transactions that would create the liquidation opportunity (referred to as forced liquidation) and then back-runs the transaction to acquire the collateral [35], 3) a searcher performs a forced liquidation, by front-running a transaction to create a liquidation opportunity and then back-running another transaction to claim the collateral. Furthermore, a collusion of searchers could manipulate market prices to extract MEV from multiple blocks by performing liquidation transactions [36]. These types of liquidation transactions have a detrimental impact on the borrowers and the market as it increases collateral requirements. Figure 17 illustrates an example of a liquidation transaction. Liquidations are of two types: fixed spread-based and auction-based [35]
@@ 5.6 Time-bandit
The previously described MEV transactions profit from re-arranging, inserting, or removing transactions in a block that is to be mined. On the contrary, time-bandit MEV transactions aim to extract MEV from already mined block [13]. The searcher, on spotting an MEV opportunity in a finalized block, would fork the blockchain by re-mining a block with MEV-acquiring transactions. Time-bandit MEV leads to blockchain instability, as the re-mining of blocks creates competing chains, which can disrupt the finality of transactions and introduce the risk of double-spending. Furthermore, time-bandit can span across domains, enabling a searcher to re-mine blocks in multiple domains, which further destabilizes the network and complicates transaction finality [31].
Having classified different types of MEV transactions, it is crucial to examine how they can be detected on-chain. Each MEV type exhibits distinct patterns, ordering, and exploitation of victim transactions, which influence the choice of detection approaches. In the following section, we explore various MEV detection approaches and assess their effectiveness in identifying different MEV types.
@@ 6.1 Front-running Detection
Torres et al. [4] focused on the detection of displacement and suppression front-running transactions101010https://github.com/christoftorres/Frontrunner-Jones, accessed on 25 February 2025. The displacement detection approach begins by identifying the front-running () and victim () transaction pairs using a sliding window that iterates over each transaction within blocks included in the window. Each transaction input bytes is divided into 4-byte n-grams. If 95%111111100% matching criteria is not considered as might be encapsulated within along with some metadata. of n-grams in a transaction corresponds to a previous transaction, a match is reported. For a matched transaction pair, the proposed heuristic then determines if is a displacement transaction if and have different senders as well as receivers (different receivers ensure that bot contracts perform displacement transactions), the gas price of the is larger than that of , and ratio between byte sequences of and is at least 25%. The proposed approach is only limited to displacement transactions performed by bot contracts. Furthermore, it may fail to detect displacement if and do not occur within the same sliding window, leading to false negatives.
@@ 6.2 Sandwich Detection
Torres et al. [4] proposed a heuristic to detect single DEX and cross-DEX sandwich transactions. The heuristic begins by identifying all possible transfer events121212A transfer event (containing sender’s address, receiver’s address, and traded amount) is emitted whenever a token is traded. within a block. The heuristic flags the presence of sandwich transactions if three events (i.e., front-run (), victim (), and back-run ()) satisfy the following conditions: Sender of and is the receiver of , and receiver of is the sender of , difference between token value of and of is less than or equal to 1%, token contract addresses for three events are same, transaction hashes for the events are different, transaction index of is smaller than that of which in turn is smaller than that of , and gas price of is greater than that of which in turn is greater than or equal to that of . The same heuristic is employed by Weintraub et al. [35] to detect sandwich transactions executed via Flashbots131313 https://github.com/a-flashbot-in-the-pan, accessed on 25 February 2025. However, this heuristic fails to identify sandwich transactions where the difference between and exceeds 1%141414Front-run transaction index:7, victim transaction index:9, back-run transaction index:10 (Block 20633424).
To detect similar types of sandwich transactions, Qin et al [18] proposed a heuristic based on the following conditions: front-run (), victim (), and back-run () transactions are in the same block, and this order, should map to one and only one , swap directions of and are the same, and that of is the opposite, either and have the same sender or same receiver smart contract, and difference between token value swapped in and swapped back in is within 10% bound. However, the heuristic cannot detect sandwich attacks where the difference between and exceeds 10%151515Front-run transaction index:21, victim transaction index:22, back-run transaction index:23 (Block 20619305).
The heuristics employed by [4, 35, 18] fail to identify sandwich transactions where the indexes of , , are not consecutive161616Front-run transaction index:0, victim transaction index:2, back-run transaction index:3 (Block 20619280) or multiple sandwich transactions occur within a single block. Furthermore, they are not capable of detecting burger, dagwood, and liquidation sandwich transactions. These limitations result in a high rate of false negatives.
@@
The previously discussed heuristics for detecting sandwich transactions [4, 35, 18, 26, 28] rely on decoding smart contracts to identify function calls and transfer events. As a result, these methods are limited to smart contracts with accessible binary files for decoding. To address this gap, Park et al. [23] proposed a heuristic171717 https://github.com/etelpmoc/arbinet, accessed on 25 February 2025 for sandwich detection that focuses on token transfer data instead of contract data. The proposed heuristic first identifies the pairs of transactions ( and ) sharing the same recipient address within a single block. Then, the transaction pairs are filtered out if one of the following conditions is met: only one token is transferred in either of the transactions, if profits for all token transfers in are greater than zero, if have all trades with zero profits, if set of token contract address in and are not same, or total profit from and is less than zero. Compared to Flashhbots with a 95.15% F1-score, the proposed approach achieved an F1-score of 98.83%. However, since this approach relies on token transfer data, identifying specific function calls becomes difficult, resulting in false positives. Furthermore, the approach is unable to detect burger, dagwood, and liquidation sandwich transactions.
@@ 6.3 Arbitrage Detection
Daian et al. [13] conducted a preliminary quantification of arbitrage in Ethereum based on a manual analysis of atomic arbitrage transactions. Later, Qin et al. [18] proposed a heuristic to detect arbitrages. According to this heuristic, a transaction is considered arbitrage if the following conditions are met: the transaction contains more than one swap event, and all events are confined within a single transaction, all swap events must form a loop, with the input of one swap being the output of the previous swap, and the input of the first swap being the output of the last swap, and the input value of each swap event should be less than or equal to the output value of the previous event. This heuristic was also employed by Weintraub et al. [35] to detect arbitrage transactions through Flashbots and Flash Loans13. However, the heuristic has several limitations. It fails to identify a cyclic arbitrage transaction if the value of the input token for a swap exceeds the value of the output token from the previous swap 181818Transaction index:2 (Block 14628933). Furthermore, it cannot detect reverse-ordered cyclic, unordered cyclic, cyclic NFT, or cyclic multi-address arbitrage transactions. Additionally, because it only considers transactions that form a cycle of swap events, it is unable to identify burn and mint or set token arbitrage transactions.
@@
Later, Park et al. [23] introduced a heuristic for detecting cyclic NFT arbitrage transactions involving swaps between ERC-20 tokens and ERC-721 NFTs. They also proposed a graph neural network model17 to identify ordered cyclic, cyclic multi-address, burn and mint, and set token arbitrage transactions. The model uses features related to acquired profit, the number of tokens sent and received, involved addresses, and the number of transfers sent and received to construct a graph. The authors implemented Graph Convolutional Networks (GCN), GraphSAGE, and Graph Attention Networks (GAT). While their approach achieved high F1-scores of 99.54% and 97.59% for training and test data, respectively, compared to Flashbots’ 68.65% and 62.85%, it may suffer from a high false positive rate due to the dynamic nature of the DEX market. Furthermore, a decrease in the performance of test data compared to train data could be due to overfitting and requires further analysis. The approach fails to detect reverse-ordered and unordered cyclic arbitrage transactions.
To detect non-atomic DEX-CEX arbitrage transactions, Heimbach et al. [32] proposed a heuristic191919https://github.com/liobaheimbach/Non-Atomic-Arbitrage-in-Decentralized-Finance, accessed on 25 February 2025 based on the following conditions: the transaction executes only one swap in a DEX and does not consume more than 400,000 gas, transaction is submitted via a private pool (more details on private pools in Section 7.3), transaction either includes a direct transfer to the block producer or has a priority fee of at least 1 GWei, transaction is either the first swap executed in the desired direction within a liquidity pool or all preceding transaction must have same recipient (to ensure the desired gains before slippage changes the market conditions), and the swap involves two tokens that are also traded on CEXes. However, the heuristic is underlined by the assumptions that non-atomic DEX-CEX arbitrage transactions are not broadcasted to the public mempool to avoid front-running, and searchers pay high-priority fees or direct transfers to the block producers to ensure the inclusion of their transaction in the desired block. Furthermore, the heuristic only focuses on non-atomic transactions.
@@ 6.4 Liquidation Detection
Weintraub et al. [35] developed a method to identify liquidation transactions13 in lending and borrowing protocols via Flashbots and Flash loans by searching for LiquidationCall and LiquidateBorrow events on the Aave and Compound lending platforms, respectively. Furthermore, Qin et al. [18] extended the scope by detecting liquidation events across Aave, Compound, and dYdX platforms. However, these approaches are limited to specific platforms.
@@ 6.5 Multi-type MEV Detection
Varun et al. [37] proposed a neural network-based pre-chain transaction screening approach to detect front-running and sandwich transactions before they are appended to the blockchain. Their approach specifically targets displacement and suppression front-running and single and cross-DEX sandwich transactions. Their approach involves extracting features for a transaction based on prior transactions and inputting them into a pre-trained Multi-Layer Perceptron (MLP) model to identify MEV transactions. The features extracted include the transaction’s gas price, the mean and standard deviation of transactions’ gas prices over the last 10 blocks, the mean and standard deviation of the same account’s historical gas prices, gas token usage, and a predicted gas price. The predicted gas price is obtained using a Long Short-Term Memory (LSTM) model based on the gas prices of the previous 12 non-MEV transactions. Using the MEV transactions data identified by Torres et al. [4] for training and validation, the approach achieved accuracies of 85.53% for suppression front-running transactions, 88.58% for displacement front-running transactions, and 89.08% for sandwich transactions. However, the proposed learning-based model may be prone to overfitting as it depends on a tailored dataset and might fail to generalize to a wider range of front-running and sandwich transactions.
@@
Li et al. [24] introduced a learning-based approach to detect sandwich, arbitrage, and liquidation MEV transactions. The authors proposed ACTLIFTER, which identifies MEV transactions within Flashbot bundles by mapping them to asset transfer patterns. Additionally, they developed ACTCLUSTER to discover new categories of MEV transactions within the sandwich, arbitrage, and liquidation types. ACTCLUSTER clusters MEV actions identified by ACTLIFTER by converting the bundle into a matrix where each DeFi action is represented as an action block. This matrix is processed using a Convolutional Neural Network (CNN) to extract features, which are then flattened and fed into three fully connected neural network layers. MEV actions are initially labeled as a sandwich, cyclic arbitrage, or liquidation transaction, with these labels being extended after each clustering round using DBSCAN. ACTLIFTER achieved better detection results compared to Etherscan202020https://etherscan.io/, accessed on 05 September 2024 and DeFiRanger [38]. However, ACTLIFTER is limited to identifying swap activities involving a token pair and fails to detect token chains.
@@
Table II presents the capabilities of the proposed detection approaches in detecting different types of MEV transactions. In the context of front-running MEV transactions, no approach focuses on replacement transactions. Regarding sandwich MEV transactions, most works focus on detecting single and cross-DEX sandwich transactions [4, 35, 18, 26, 23, 37, 24]. Only a few focus on detection of burger [26, 24], dagwood [26], and liquidation sandwich [28, 24] transactions. No single approach is capable of detecting all types of sandwich transactions. Furthermore, all these approaches are limited to single-block sandwich transactions. Consequently, the detected sandwich transactions represent only lower bounds and could result in a high number of false negatives [4, 18]. For arbitrage detection, most works focus on ordered cyclic transactions [35, 18, 26, 23, 30, 24]. Additionally, most of the works [35, 18, 26, 23, 24] are limited to atomic arbitrage transactions, where multiple swap events occur within a single transaction. Only [30] addresses arbitrages performed through both atomic and non-atomic transactions. Lastly, [35, 18, 24] focus on the detection of liquidation MEV transactions.
Many detection approaches can be combined to improve MEV transaction identification. Integrating the rules from heuristic-based detection approaches [4, 35, 18, 26, 28, 23, 30], enables the detection of multiple types of MEV transactions. However, the combined heuristic-based approach, without introducing additional rules, fails to detect replacement, back-running, multi-block sandwich, multi-address arbitrage, burn and mint arbitrage, set token arbitrage, and time bandit MEV transactions. Similarly, while combining learning-based detection approaches [37, 24, 23] enhances MEV detection capabilities, it still cannot detect replacement, back-running, dagwood, multi-block sandwich, unordered cyclic arbitrage, reverse ordered cyclic arbitrage and time bandit MEV transactions. In addition, integrating learning-based approaches is challenging due to differences in feature extraction and detection methodologies.
@@
To summarize, no combined approach is capable of detecting all types of MEV transactions. Furthermore, the detection results presented by the studies discussed above [4, 35, 18, 26, 28, 23, 13, 30, 37, 24] are only estimates of MEV, as there is no definitive ground truth for MEV [18]. In addition, the works employing heuristic methods [4, 35, 18, 26, 28, 23, 30] represent reactive MEV detection approaches, focusing on quantifying MEV acquired in the past. Estimating potential or future MEV via a proactive approach is challenging because the MEV obtained by each searcher varies based on the optimization techniques used to maximize MEV and the targeted victim transactions [19].
While MEV detection approaches help identify and analyze MEV transactions, managing MEV exploitation requires more than just identification. MEV mitigation strategies aim to proactively minimize or democratize MEV exploitation. In the following section, we analyze different MEV mitigation strategies and assess their effectiveness in limiting different types of MEV transactions.
| MEV transactions | |||||||||||||||||||
| Front-running | Sandwich | Arbitrage | |||||||||||||||||
@@
| Detection approach | Displa- cement | Repla- cement | Suppre- ssion | Back- running | Single DEX | Cross- DEX | Burger | Dagwood | Liquid- ation | Ordered cyclic | Reverse- ordered cyclic | Unordered cyclic | Cyclic NFT | Cyclic multi- address | Burn and mint | Set token | Cross- domain | Liquid- ation | Time- bandit |
| [4] | |||||||||||||||||||
| [35] | |||||||||||||||||||
| [18] | |||||||||||||||||||
| [26] | |||||||||||||||||||
| [28] | |||||||||||||||||||
| [23] | |||||||||||||||||||
| [30] | |||||||||||||||||||
| [37] | |||||||||||||||||||
| [24] | |||||||||||||||||||
| [32] | |||||||||||||||||||
@@
| : detected; : not detected; : partially detected; 1if and occur within the same block range window; 2if multiple suppression transactions are sent to the same bot contract; 3if difference between and is 1%; 4if the input value of a swap event the output value of previous event; 5if difference between and is 10%; 6if and has the same swap directions; 7only DEX-CEX | |||||||||||||||||||